# Flexible Physical Camouflage Generation Based on a Differential Approach

### v1

@@ 3.4.1 Adversarial Loss
3.4.2 Texture Smoothness
Texture smoothness is a key procedure to ensure the generated texture can be applied in the physical world with naturalness, effectiveness, and printability. As a smooth texture is more natural and realistic, making it less suspicious to human observers and more difficult to detect by deep learning systems, also improves the effectiveness of the adversarial attack by reducing the visual artifacts and noise that may interfere with the target object recognition, and is more printable and transferable to physical objects, making it easier to apply the adversarial attack in the real world. Therefore, we propose the total variation loss in Equation 5.
$L_{TV}=\frac{1}{N_{nz}}\sum_{i,j}(\|T_{i,j}-T_{i+1,j}\|)+(\|T_{i,j}-T_{i,j+1}),$ (5)
@@ 3.4.2 Texture Smoothness
3.4.3 Digital to Physical Score
It is necessary to convert the digital color to the physical world to ensure the generated texture can be applied in the physical world. The generally used method is color management to ensure the color in digital and printable spaces is the same. Let $P$ be the set of printable colors, and let $C$ be the set of colors in the adversarial texture. Let $d(T_{i},p_{j})$ be the Euclidean distance between color $T_{i}$ in the texture and color $p_{j}$ in the set of printable colors. Then, the non-print-ability score loss $L_{NPS}$ is given by:
@@ 3.4.3 Digital to Physical Score
3.4.4 Concealment Constraint
We propose a concealment constraint loss to ensure the similarity between the generated texture and the background, enhancing the dual concealment of texture in both machine vision and human visual observation. First, we compute the average RGB value ${c_{r}}$ within the background region based on the target. Based on the set color fluctuation threshold $c_{r_{u}}$, the range of background color variation can be obtained as $[u_{l},u_{h}]$, where $u_{l}=c_{r}-c_{r_{u}}$ and $u_{h}=c_{r}+c_{r_{u}}$. Then, we compute the difference between the generated texture and the background color variation to construct the concealment constraint loss, which can be represented as Equation 7.
$L_{CR}=\frac{1}{M}\sum_{i,j}\|T_{i,j}-u_{l}\|+\|T_{i,j}-u_{h}\|,$ (7)
where $M$ denotes the total number of the nonzero pixels, $T_{i,j}$ represents the pixel value of adversarial texture at coordinate $(i,j)$, lower $L_{CR}$ indicates closer resemblance between the texture color and the background. Thus, minimizing $L_{CR}$ enhances the similarity between texture color and background, thereby improving concealment in the environment.
@@ 3.4.4 Concealment Constraint
2:  Obtain the input image $x$, the ground truth $y$
@@
5:        Obtain the adversarial texture $T_{k}$ with Equation 1
6:        Obtain the rendered image $x_{k}$ with Algorithm 1
7:     end for
8:     $b^{(x)},c^{(x)}=\mathbf{O}(x_{k},y^{(x)})$
9:     Obtain the adversarial loss with Equation 4
10:     Obtain the smooth variants with Equation 5
11:     Obtain the color constraints with Equation 6
12:     Obtain the concealment constraints with Equation 7
13:     Set $L$ by Equation 8
14:     Update diffusion model parameters $\theta$ for minimizing $L$ by via backpropagation
15:  end for
4 Experiments
@@ 4.1 Experimental Settings
Implementation Details In the actual deployment of adversarial camouflage, there is a process of magnification and printing. To minimize the loss of details in adversarial camouflage, considering the existing hardware conditions, we set a higher resolution for the adversarial camouflage, with an initial size of 480*480. The adversarial camouflage uses the same random seed for each training iteration, initialized with random noise. We employ Adam as the optimizer with default parameters, a learning rate of 0.001, and a maximum of 10 epochs. The parameters of the loss function are configured based on extensive experimental experience, with the following settings: $\alpha=0.05$, $\beta=1.0$, $\gamma=1.0$, $\mu=2.5$, $\tau=2.0$. All our code is implemented in PyTorch and training/validation are conducted on an NVIDIA GeForce RTX 3090 GPU.
4.2 Universality Evaluations
@@
Transferability Evaluation Across Different Victim Models We trained adversarial camouflages against different victim models: FPA(v3) against YOLOv3, FPA(v5) against YOLOv5, and FPA(FR) against FRCNN.To ensure a fair comparison with prior adversarial camouflage works, we validated our approach using the same dataset as FCA. We added 5 groups of adversarial camouflages as comparative experiments: Noise, CAMOU[57], ER[53], DAS[48], FCA[47]. Among these, CAMOU and ER adversarial camouflage cover the entire vehicle by repeated pasting. For DAS and FCA adversarial camouflage, we employed the same neural rendering technique as the original paper, overlaying textures onto the vehicle surface through neural rendering. As shown in Figure.3, vehicles covered by our trained adversarial camouflage interfere with detector recognition and demonstrate the most effective attack. Table.1 presents the comparison results of adversarial camouflage on AP@0.5. Our method achieves the best results across multiple detection models. Specifically, on YOLOv3, FPA(v5) exhibits the maximum drop of 73.9% in AP@0.5, surpassing FCA by 10%. On the widely used YOLOv5, it achieves an 89.1% reduction, indicating the concealment capability of our camouflage at any angle. Due to the architectural differences of FRCNN and the diversity of training datasets for detection models, the effectiveness of FPA (v5) in attacks slightly decreased, yet it still achieved a 52% reduction, surpassing other real-world attack methods. When employing FRCNN as the victim model, our adversarial camouflage, FPA(FR), achieved a reduction of 74.9%. The last column in the table indicates the average decrease in AP@0.5 across the three detectors, with our method achieving the best results. Particularly, FPA(v5) achieved a reduction of 71.6%, demonstrating the ability of our framework to generate highly aggressive camouflage when facing different victim detection models.
@@
Transferability Evaluation Across Different Architecture Detectors To assess the effectiveness of adversarial camouflage on unknown models, we introduced several different architectures of detection models, including SSD, YOLOv7, Mask R-CNN(MkRCNN), Cascade R-CNN(CaRCNN), DETR, and RT-DETR. The test results for AP@0.5 are shown in Table.2, where our adversarial camouflage, FPA(v5), outperforms others on every detection model. In the latest detection model YOLOv7, it achieves a significant reduction from the original 97.9% to 11.2%, demonstrating robust attack effectiveness. When confronted with CaRCNN, while other camouflage methods only slightly decrease AP@0.5, our adversarial camouflage, FPA(v5), achieves a reduction of 48.2%. Interestingly, when facing detection models based on Transformer architecture, our attack method also demonstrates strong attack capability, achieving a reduction of 68.3% in DETR. This indicates the robustness of our method against unknown architectures of black-box models. Figure.4 illustrates the prediction results of our adversarial camouflage, FPA(v5), across different detection models. Our method enables vehicles to evade detection model recognition, further validating the robust transferability of our approach.
Table 2: The AP@0.5 test results of adversarial camouflage under different detection models.
@@
Transferability Evaluation Across Different Tasks We aim for the adversarial camouflage, once covering the vehicle, to not only impact target detection recognition but also disrupt other tasks, such as depth estimation commonly used in the field of autonomous driving. We validate the attack performance of the adversarial camouflage on publicly trained depth estimation models, including Monodepth2 (MoD for short)[14], Depth Hints (DH for short)[51], and ManyDepth (MaD for short)[52]. For a more comprehensive evaluation, we generated 24 test images based on different environments at distances of 5 meters and 10 meters. Additionally, we selected two common metrics on depth estimation, depth error$({\mathcal{E}}_{d})$, and the proportion of affected regions$(\mathcal{R}_{a})$[7]. The experimental results, as presented in Table.3, indicate that our attack achieves the best performance on MoD, with a depth influence of 12.36 and a proportion of affected regions reaching 85.4%. Figure.5 illustrates the prediction results on MoD. After applying the adversarial camouflage, the display of the original vehicle depth almost completely disappears, further confirming the strong transferability of our method across different tasks.
Table 3: Results of depth error and the proportion of affected regions testing on trained depth estimation models after applying camouflage coverage.
@@
4.3 Simulation Evaluation of Physical World Factors
@@
Robustness Evaluation under Different Occlusions In the physical world, targets often appear partially visible in images, leading to partial camouflage coverage. Therefore, we conducted experiments to evaluate the robustness of our method under different degrees of occlusion. To prevent images from becoming too small after occlusion, we set the distance range to [5, 10, 15, 20], pitch angles to [15, 30, 45, 60], and collected images at intervals of 12° within a 360° angle range. To generate rendered images with different occlusion levels, we start with the width of the target as the initial length and occlude it from left to right in 10 equal proportions from 0 to 1. Meanwhile, we designate proportions from 0.1 to 0.3 as small occlusion, proportions from 0.4 to 0.6 as middle occlusion, and proportions from 0.7 to 0.9 as large occlusion. Hence, each data point in Table.4 represents the ASR calculated across 360 test images. In Figure.7, we present the prediction results under occlusion ratios of [0.1, 0.4, 0.7]. From the data in the table, it is evident that our camouflage achieves an ASR close to 100% under different occlusion areas, indicating the strong robustness of our method against occlusion.
Table 4: The ASR test results of the adversarial camouflage under different occlusion ranges.
@@
Robustness Evaluation under Different Reflection Coefficients The intensity of lighting in the physical world can significantly affect adversarial camouflage. This is primarily because varying light intensities can result in different degrees of reflection on the surface of the vehicle, consequently altering the camouflage. Our method simulates physical parameters during the training process of adversarial camouflage using PyTorch3D, ensuring robustness in real-world environments. Therefore, in this section of the experiment, we used PyTorch3D to set the reflectivity coefficient(diffuse_color) of the target surface, simulating the reflection of the target under different conditions. Specifically, we defined a set of reflectivity coefficients [1.0, 3.0, 5.0, 10.0, 15.0] for experimentation. The data collection method was the same as the occlusion experiment. Each value in Table.5 is calculated based on ASR from 480 test images. From the data in the table, we can find that the ASR of adversarial camouflage fluctuates around 97% and is very little affected by reflection. The graph depicted in Figure.8 illustrates that the camouflaged vehicle successfully evades detection by the target detection system, indicating that our method exhibits good robustness across different reflectivity coefficients.
Table 5: The ASR test results of the adversarial camouflage under different reflective coefficient settings.
diffuse_color 1 3 5 10 15
@@
Figure 8: Rendered images (from left to right) at reflectance factors of 1.0, 5.0, and 15.0, and predicted results under YOLOv5.
4.4 Concealment Evaluation
@@
In all three scenarios, the adversarial camouflage generated by Camou-color achieved better results in both SSIM and CSIM. Table.6 presents the ASR results of adversarial camouflage under different training modes in YOLOv5, where “None" indicates the detection results of the vehicle without camouflage. The attack effectiveness of adversarial camouflage under the Camou-color training mode is slightly lower compared to Camou-raw, but both can significantly impact the target detection model. Figure.11 illustrates rendered images of vehicles concealed with camouflage across different scenarios. Visually, these images create confusion, indicating that our method significantly enhances camouflage concealment while maintaining attack effectiveness.
Table 6: The original vehicle, along with the ASR test outcomes of the adversarial camouflage in various scenarios obtained through the two training methodologies.
@@
4.5 Physical World Evaluation
@@
In this section, we evaluated the adversarial camouflage’s performance in the physical world. Specifically, we conducted experiments in the physical world using a 1:24 scale Audi Q5 car model. Due to budget and time constraints, we utilized an HP color printer to print our adversarial camouflage FPA(v5), and then cut and pasted the camouflage parts onto the corresponding areas of the car model. We captured surround shots of the target at three different distances within three distinct environments to assess the efficacy of camouflage, utilizing the Redmi K50 Pro as the imaging device. The original and camouflaged cars were captured using the same model for data collection. From the collected data, we extracted 200 images for validation purposes, with the AP@0.5 results presented in Table.7, and a selection of prediction outcomes illustrated in Figure.12. Our adversarial camouflage significantly reduces AP@0.5 across various detection models, particularly in SSD, showing a reduction of up to 82.5%. This indicates that our method can effectively overcome physical world interference and efficiently attack object detection models.
Table 7: The AP@0.5 test results of the original car and the camouflaged car in the physical world.
Method SSD Yolov5 FRCNN MkRCNN DETR RT-DETR
RAW 0.930 0.995 0.995 0.989 0.985 0.977
FPA(v5) 0.105 0.451 0.529 0.475 0.251 0.488
@@
Figure 12: The prediction results of the original model car and the camouflaged car in the physical world.
4.6 Ablation Studies

### v3

@@ III-D Loss constraints
Texture Smoothness. Texture smoothness is a key procedure to ensure the generated texture can be applied in the physical world with naturalness, effectiveness, and printability. As a smooth texture is more natural and realistic, making it less suspicious to human observers and more difficult to detect by deep learning systems, also improves the effectiveness of the adversarial attack by reducing the visual artifacts and noise that may interfere with the target object recognition, and is more printable and transferable to physical objects, making it easier to apply the adversarial attack in the real world. Therefore, we propose the total variation loss in Equation 8.
$L_{TV}=\frac{1}{N_{nz}}\sum_{i,j}(T_{i,j}-T_{i+1,j})^{2}+(T_{i,j}-T_{i,j+1})^{2},$ (8)
@@
Digital to Physical Score. It is necessary to convert the digital color to the physical world to ensure the generated texture can be applied in the physical world. The generally used method is color management to ensure the color in digital and printable spaces is the same. Let $P$ be the set of printable colors, and let $C$ be the set of colors in the adversarial texture. Let $d(T_{i},p_{j})$ be the Euclidean distance between color $T_{i}$ in the texture and color $p_{j}$ in the set of printable colors. Then, the non-print-ability score loss $L_{NPS}$ is given by:
@@
Concealment Constraint. We propose a concealment constraint loss to ensure the similarity between the generated texture and the background, enhancing the dual concealment of texture in both machine vision and human visual observation. First, we compute the average RGB value ${c_{r}}$ within the background region based on the target. Based on the set color fluctuation threshold $c_{r_{u}}$, the range of background color variation can be obtained as $[u_{l},u_{h}]$, where $u_{l}=c_{r}-c_{r_{u}}$ and $u_{h}=c_{r}+c_{r_{u}}$. Then, we compute the difference between the generated texture and the background color variation to construct the concealment constraint loss, which can be represented as Equation 10.
$L_{CR}=\frac{1}{N}\sum_{i,j}(T_{i,j}-u_{l})^{2}+(T_{i,j}-u_{h})^{2},$ (10)
where $N$ denotes the total number of the nonzero pixels, $T_{i,j}$ represents the pixel value of adversarial texture at coordinate $(i,j)$, lower $L_{CR}$ indicates closer resemblance between the texture color and the background. Thus, minimizing $L_{CR}$ enhances the similarity between texture color and background, thereby improving concealment in the environment.
@@
2:  Load the background images $x_{bk}$ and the UV map $T_{UV}$
@@
5:        Obtain the adversarial texture $T_{k-1}$ with Equation 3
6:     end for
7:     Obtain the desired adversarial texture $T_{adv}$ from the generated texture $T_{k-1}$ with $T_{adv}=T_{k-1}\cdot T_{UV}$.
8:     Obtain the rendered image $x_{ren}$ and the mask image $x_{mask}$ with Algorithm 1.
9:     Obtain the true label $y$ corresponding to $x_{ren}$ by identifying the boundaries of the target region in $x_{mask}$.
10:     Obtain the adversarial example $x$ with Equation 6.
11:     <$b^{(x)},c^{(x)}>=\mathbf{O}(x,y^{(x)})$
12:     Obtain the adversarial loss with Equation 7.
13:     Obtain the smooth variants with Equation 8.
14:     Obtain the color constraints with Equation 9.
15:     Obtain the concealment constraints with Equation 10.
16:     Set $L$ by Equation 11.
17:     Update diffusion model parameters $\theta$ for minimizing $L$ by via backpropagation.
18:  end for
IV Experiments
@@ IV-A Experimental Settings
Implementation Details In the actual deployment of adversarial camouflage, there is a process of magnification and printing. To minimize the loss of details in adversarial camouflage, considering the existing hardware conditions, we set a higher resolution for the adversarial camouflage, with an initial size of 480*480. The adversarial camouflage uses the same random seed for each training iteration, initialized with random noise. We employ Adam as the optimizer with default parameters, a learning rate of 0.001, and a maximum of 10 epochs. We set $\upsilon_{model}$ to define the lighting mode: in the range [0, 0.33] for point light mode, in the range (0.33, 0.67] for directional light mode, and in the range (0.67, 1] for global illumination mode. The range for the light intensity parameter $\upsilon_{light}$ is set to [0.2, 5], and the range for the material reflectivity $\upsilon_{material}$ is set to [0.5, 1.5]. The parameters of the loss function are configured based on extensive experimental experience, with the following settings: $\alpha=0.05$, $\beta=1.0$, $\gamma=1.0$, $\mu=2.5$, $\tau=2.0$. All our code is implemented in PyTorch and training/validation are conducted on an NVIDIA GeForce RTX 3090 GPU.
IV-B Universality Evaluations
Transferability Evaluation Across Different Victim Models We trained adversarial camouflages against different victim models: FPA(v3) against YOLOv3, FPA(v5) against YOLOv5, and FPA(FR) against FRCNN.To ensure a fair comparison with previous adversarial camouflage works, we validated our approach using the same dataset as FCA. We added eight groups of adversarial camouflages as comparative experiments: Noise, CAMOU[60], ER[56], DAS[51], FCA[50], DTA[44], ACTIVE[45], RAUCA[63]. Among these, the adversarial camouflage CAMOU and ER cover the entire vehicle by repeated pasting. For DAS, FCA and RAUCA adversarial camouflage, we employed the same neural rendering technique as the original paper, overlaying textures onto the vehicle surface through neural rendering. For DTA and ACTIVE, we applied the repeated texture projection method to map them onto the target surface. To validate the effectiveness of the FPA framework, we re-generated adversarial textures FPA(bk) for YOLOv5 using 20 background images collected from Carla (excluding cars).
@@ IV-B Universality Evaluations
As shown in Figure.4, vehicles covered by our trained adversarial camouflage interfere with detector recognition and demonstrate the most effective attack. Table.I presents the comparison results of adversarial camouflage on AP@0.5. Our method achieves the best results on multiple detection models. Specifically, on YOLOv3, FPA(v5) exhibits the maximum drop of 73.9% in AP@0.5, surpassing FCA by 10%. On the widely used YOLOv5, it achieves an 89.1% reduction, indicating the concealment capability of our camouflage at any angle. Due to the architectural differences of FRCNN and the diversity of training datasets for detection models, the effectiveness of FPA (v5) in attacks slightly decreased, yet it still achieved a 52% reduction, surpassing other real-world attack methods. When employing FRCNN as the victim model, our adversarial camouflage, FPA(FR), achieved a reduction of 74.9%. The last column in the table indicates the average decrease in AP@0.5 across the three detectors, with our method achieving the best results. In particular, FPA (v5) achieved a reduction of 71.6%, demonstrating the ability of our framework to generate highly aggressive camouflage when faced with different victim detection models. The validation results of FPA(bk) are only slightly lower than those of FPA(v5), yet it still demonstrates strong attack capabilities. This shows that our framework can generate adversarial textures from a small number of background images, and the adversarial textures exhibit strong robustness against different detectors and viewpoint changes.
@@
Transferability Evaluation Across Different Architecture Detectors To assess the effectiveness of adversarial camouflage on unknown models, we introduced several different architectures of detection models, including SSD, YOLOv7, Mask R-CNN(MkRCNN), Cascade R-CNN(CaRCNN), DETR, and RT-DETR. The test results for AP@0.5 are shown in Table.II, where our adversarial camouflage, FPA(v5), outperforms others on every detection model. In the latest detection model YOLOv7, it achieves a significant reduction from the original 97.9% to 11.2%, demonstrating robust attack effectiveness. When confronted with CaRCNN, while other camouflage methods only slightly decrease AP@0.5, our adversarial camouflage, FPA(v5), achieves a reduction of 48.2%. Interestingly, when facing detection models based on Transformer architecture, our attack method also demonstrates strong attack capability, achieving a reduction of 68.3% in DETR. This indicates the robustness of our method against unknown architectures of black-box models. From the table, we can see that the overall attack effect of RAUCA is slightly lower than that of FPA. However, RAUCA is trained based on a constructed weather dataset (40,960 images). In order to achieve anti-interference of environmental factors, an environmental feature extractor is added to combine with Neural Render. Compared with the method in this paper, the structure is more complex, the amount of data required is larger, the deployment is more cumbersome, and the attack effect is slightly worse. Figure.5 illustrates the prediction results of our adversarial camouflage, FPA(v5), across different detection models. Our method enables vehicles to evade detection model recognition, further validating the robust transferability of our approach.
TABLE II: The AP@0.5 test results of adversarial camouflage under different detection models.
@@
Transferability Evaluation Across Different Tasks We aim for the adversarial camouflage, once covering the vehicle, to not only impact target detection recognition but also disrupt other tasks, such as depth estimation commonly used in the field of autonomous driving. We validate the attack performance of the adversarial camouflage on publicly trained depth estimation models, including Monodepth2 (MoD for short)[15], Depth Hints (DH for short)[54], and ManyDepth (MaD for short)[55]. For a more comprehensive evaluation, we generated 24 test images based on different environments at distances of 5 meters and 10 meters. Additionally, we selected two common metrics on depth estimation, depth error$({\mathcal{E}}_{d})$, and the proportion of affected regions$(\mathcal{R}_{a})$[8]. The experimental results, as presented in Table.III, indicate that our attack achieves the best performance on MoD, with a depth influence of 12.36 and a proportion of affected regions reaching 85.4%. Figure.6 illustrates the prediction results on MoD. After applying the adversarial camouflage, the display of the original vehicle depth almost completely disappears, further confirming the strong transferability of our method across different tasks.
TABLE III: Results of depth error and the proportion of affected regions testing on trained depth estimation models after applying camouflage coverage.
@@
IV-C Simulation Evaluation of Physical World Factors
@@
Robustness Evaluation under Different Occlusions In the physical world, targets often appear partially visible in images, leading to partial camouflage coverage. Therefore, we conducted experiments to evaluate the robustness of our method under different degrees of occlusion. To prevent images from becoming too small after occlusion, we set the distance range to [5, 10, 15, 20], pitch angles to [15, 30, 45, 60], and collected images at intervals of 12° within a 360° angle range. To generate rendered images with different occlusion levels, we start with the width of the target as the initial length and occlude it from left to right in 10 equal proportions from 0 to 1. Meanwhile, we designate proportions from 0.1 to 0.3 as small occlusion, proportions from 0.4 to 0.6 as middle occlusion, and proportions from 0.7 to 0.9 as large occlusion. Hence, each data point in Table.IV represents the ASR calculated across 360 test images. In Figure.8, we present the prediction results under occlusion ratios of [0.1, 0.4, 0.7]. From the data in the table, it is evident that our camouflage achieves an ASR close to 100% under different occlusion areas, indicating the strong robustness of our method against occlusion.
TABLE IV: The ASR test results of the adversarial camouflage under different occlusion ranges.
@@
Robustness Evaluation under Different Reflection Coefficients The intensity of lighting in the physical world can significantly affect adversarial camouflage. This is primarily because varying illumination intensities can result in different degrees of reflection on the surface of the vehicle, consequently altering the camouflage. Our method simulates physical parameters during the training process of adversarial camouflage using PyTorch3D, ensuring robustness in real-world environments. Therefore, in this part of the experiment, we used PyTorch3D to set the environmental illumination intensity ($\upsilon_{light}$) to simulate the changes in the target surface under different conditions. Specifically, we defined a set of illumination intensity coefficients [0.5, 1.0, 2.0, 3.0, 5.0] for the experiments. The data collection method was the same as in the occlusion experiments. Each value in Table V is calculated based on the ASR from 480 test images. From the data in the table, it can be seen that the ASR of the adversarial camouflage fluctuates around 97%, showing little to no impact from changes in illumination. The graphs depicted in Figure 9 illustrate that the camouflaged vehicles successfully evade detection by the target detection system, indicating that our method exhibits strong robustness under varying lighting conditions.
TABLE V: The ASR test results of the adversarial camouflage under different illumination intensity settings.
diffuse_color 0.5 1.0 2.0 3.0 5.0
@@
Figure 9: The rendered images with illumination intensities of 0.5, 2.0, and 5.0 (from left to right) and the corresponding prediction results under YOLOv5 are shown.
IV-D Concealment Evaluation
@@
In all three scenarios, the adversarial camouflage generated by Camou-color achieved better results in both SSIM and CSIM. Table.VI presents the ASR results of adversarial camouflage under different training modes in YOLOv5, where “None" indicates the detection results of the vehicle without camouflage. The attack effectiveness of adversarial camouflage under the Camou-color training mode is slightly lower compared to Camou-raw, but both can significantly impact the target detection model. Figure.12 illustrates rendered images of vehicles concealed with camouflage across different scenarios. Visually, these images create confusion, indicating that our method significantly enhances camouflage concealment while maintaining attack effectiveness.
TABLE VI: The original vehicle, along with the ASR test outcomes of the adversarial camouflage in various scenarios obtained through the two training methodologies.
@@
IV-E Physical World Evaluation
@@
In this section, we evaluated the adversarial camouflage’s performance in the physical world. Specifically, we conducted experiments in the physical world using a 1:24 scale Audi Q5 car model. Due to budget and time constraints, we utilized an HP color printer to print our adversarial camouflage FPA(v5), along with four comparative camouflages: RAUCA, ACTIVE, DTA, and FCA. We then cut out the camouflaged sections and adhered them to the corresponding areas of the car models. We captured surround shots of the target at three different distances within three distinct environments to assess the efficacy of camouflage, utilizing the Redmi K50 Pro as the imaging device. The original and camouflaged cars were captured using the same model for data collection. We collected 100 images for each type of camouflage to validate the results. The AP@0.5 results are shown in Table.VII, and some of the prediction results can be seen in Figure.13. Our adversarial camouflage significantly reduces AP@0.5 across different detection models. Notably, it achieves a reduction of up to 66.7% in the SSD model. While other adversarial camouflages struggle to generate effective attacks against the two-stage detector Faster R-CNN, our camouflage successfully lowers AP@0.5 to 0.529. This demonstrates that our method has stronger attack effects in the physical world than previous work.
TABLE VII: The AP@0.5 test results of the original car and the camouflaged car in the physical world.
Method SSD Yolov5 FRCNN DETR
RAW 0.930 0.995 0.995 0.985
RAUCA 0.395 0.586 0.758 0.406
ACTIVE 0.571 0.675 0.891 0.607
DTA 0.854 0.875 0.912 0.725
FCA 0.713 0.862 0.874 0.621
FPA 0.263 0.442 0.529 0.315
@@
Figure 13: The prediction results of the original model car and the camouflaged cars in the physical world. Green boxes indicate correct detections by the detector, while red boxes indicate successful attacks by the adversarial camouflage.
IV-F Ablation Studies